# DRAM Errors and Cosmic Rays: Space Invaders or Science Fiction?

### v2

@@ Abstract
Our study describes the underlying sources of high-energy cosmic rays, explains which particles should be observed, and points to the publicly-available logs from reliable measurement facilities (Section II). We study the DRAM errors of MareNostrum 3 large-scale production system and mid-scale Mont-Blanc prototype, which does not support memory error correction in hardware. The systems are located in the same geographical location, in Barcelona (Spain), and the errors are logged in overlapping time-periods. The MareNostum 3 logs cover 2000 billion MB-hours of server-class DDR3-1600 DIMMs during which we detected 4.5 million corrected and 71 uncorrected DRAM errors (Section III). In the Mont-Blanc prototype, we collected over 25 million memory errors during the 135 million MB-hours production of mobile LPDDR3-1600 chips (Section IV).
@@ II-B Data collection
Because of proximity, we used the Guadalajara cosmic ray logs. To ensure that the measurements accurately reflect cosmic ray activity, we use data corrected for pressure and instrumental factors.111In the Neutron Monitor logs, this is referred to as efficiency correction. Corrections for instrumental changes, such as modifications of the building and the measurement equipment, are necessary to maintain data consistency. Also, the atmospheric pressure correction is important because higher atmospheric pressure increases the air mass above the detector, reducing the detected neutron flux.
@@ V-B DRAM errors
Figure 2(a) shows the MareNostrum corrected errors (CEs) hourly and monthly occurrences between October 2014 and November 2016. On most hours there are up to 100 corrected errors, but on a few hours the number of errors is large, up to about 10,000. The results suggest that we should further explore a potential alignment between the neutron count and error peaks. Comparison of the neutrons and CEs monthly trend-lines, Figures 1 and 2(a), detects no obvious correlation between these two variables. We repeat this analysis focused only on the transient CEs, and we reach the same conclusions.
@@ IX The signal and the noise
For comparison, we also plot the hour-of-the-day for MareNostrum errors between February 2015 and February 2016. In this period, we detected 29 uncorrected errors (Figure 8(c)). We also plot the hour-of-the-day chart for all corrected errors (Figure 8(d)) and the transient ones (Figure 8(e)). For all CEs, we detect a clear peak at 9 h, while the largest number of transient CEs is detected at 18 h. To illustrate how a small number of DIMMs can change the general trend of the data, we remove 1% of the DIMMs with the highest number of errors. These results are shown with darker narrow bars in Figures 8(d) and 8(e). We still detect some variation between the time of the day, but some of the highest error rates are now recorded in different hours. We repeated the same analysis with our full logs covering the period October 2014 to November 2016. These results are not displayed, but the conclusions are the same.

### v3

@@ Abstract
Our study describes the underlying sources of high-energy cosmic rays, explains which particles should be observed, and points to the publicly-available logs from reliable measurement facilities (Section II). We study the DRAM errors of MareNostrum 3 large-scale production system and mid-scale Mont-Blanc prototype, which does not support memory error correction in hardware. The systems are located in the same geographical location, in Barcelona (Spain), and the errors are logged in overlapping time-periods. The MareNostum 3 logs cover 2000 billion MB-hours of server-class DDR3-1600 DIMMs during which we detected 4.5 million corrected and 71 uncorrected DRAM errors (Section III).111Megabyte-hours (MB-hours) combines both the amount of memory used and the duration for which it is used. Specifically, it quantifies how much memory (in megabytes) is consumed over a specific period of time (in hours). In the Mont-Blanc prototype, we collected 25 million errors during the 135 million MB-hours production of mobile LPDDR3 memory (Section IV).
@@ II-B Data collection
Because of proximity, we used the Guadalajara cosmic ray logs. To ensure that the measurements accurately reflect cosmic ray activity, we use data corrected for pressure and instrumental factors. In the Neutron Monitor logs, this is referred to as efficiency correction. Corrections for instrumental changes, such as modifications of the building and the measurement equipment, are necessary to maintain data consistency. Also, the atmospheric pressure correction is important because higher atmospheric pressure increases the air mass above the detector, reducing the detected neutron flux.
@@ V-B DRAM errors
Figure 2(a) shows the MareNostrum corrected errors (CEs) hourly and monthly occurrences between October 2014 and November 2016. On most hours there are up to 100 corrected errors, but on a few hours the number of errors is large, up to about 10,000. The results suggest that we should further explore a potential alignment between the neutron count and error peaks. Comparison of the neutrons and CEs monthly step-lines, Figures 1 and 2(a), detects no obvious correlation between these two variables. We repeat this analysis focused only on the transient CEs (not depicted in Figure 2), and we reach the same conclusions.
@@ IX The signal and the noise
For comparison, we also plot the hour-of-the-day for MareNostrum errors between February 2015 and February 2016. In this period, we detected 29 uncorrected errors (Figure 8(c)). We also plot the charts for all corrected errors (Figure 8(d)) and the transient ones (Figure 8(e)). For all CEs, we detect a clear peak at 9 h, while the largest number of transient CEs is detected at 18 h. To illustrate how a small number of DIMMs can change the general trend of the data, we remove 1% of the DIMMs with the highest number of errors. These results are shown with darker narrow bars in Figures 8(d) and 8(e). We still detect some variation between the time of the day, but some of the highest error rates are now recorded in different hours. We repeated the same analysis with our full logs covering the period October 2014 to November 2016. These results are not displayed, but the conclusions are the same.
@@ X Related Work
Physics-based error models require the simulation of physical processes, nuclear reactions and radiation transport models, that generate a charge capable of impacting the functionality of electronic circuits. This simulation requires complex commercial tools and publicly unavailable circuit design details. Purely physics-based error models are impracticable even for device manufacturers because the simulation (solving coupled differential equations) is typically beyond the reach of available computing power [51]. Additionally, DRAM error models should consider memory traffic, data values and access patterns [8, 52, 53, 54], which can vary greatly depending on the customer and their target applications.
As an alternative, semi-empirical models calibrate the projections of physics-based simulations with error rates from experiments in which devices are irradiated with neutron beams of known flux [51, 55, 20, 52]. Beam experiments, however, also present challenges. The first one is providing consistent conditions among different tests [8, 52]. Since 2001, these conditions have been standardized by JEDEC [53, 54]. Second, the complexity of these experiments limits the number of studies. In the last couple of decades only three studies have reported beam-test results for DRAM and HBM devices [56, 57, 58]. Finally, beam experiments may not match production operating conditions [8, 52]. For example, the spectrum of neutron energies and fluxes can vary significantly between different geographical locations [51, 20, 8, 52]. For this reason, recent JEDEC standards include a flux calculator to determine the relative neutron flux at a particular location and adjust it with respect to a reference location (New York City) [59]. However, this flux calculation does not account for real-time variations in cosmic-ray intensity, such as those due to local atmospheric pressure or solar and cosmic events [20]. In our study, this variation is depicted in Figure 1. Some studies report much higher variations of up to 30% over just a few days [51]. Other factors that are difficult to correct include the presence of buildings and whether the system is located on the top floor or in the basement [51].
Overall, the analysis of DRAM errors induced by cosmic rays presents many challenges. In-field studies are essential because they report device error rates under production conditions: running representative workloads and datasets within a representative hardware and software infrastructure [52, 60]. Additionally, due to the lack of detailed physical device models and beam-test results, in-field studies currently serve as the primary source of information on DRAM error rates.
@@ XI Conclusions
Acknowledgment
The research was supported by the Spanish Government through the contracts PID2019-107255GB-C21, PID2023-146511NB-I00 and CEX2021-001148-S, funded by MCIN/AEI/10.13039/501100011033. The work also received funding from the Department of Research and Universities of the Government of Catalonia (Code: 2021 SGR 00807, 2021 SGR 01007, 2021 SGR 01264). Paul Carpenter holds the Ramon y Cajal fellowship RYC2018-025628-I funded by MICIU/AEI/10.13039/501100011033 and ”ESF Investing in your future”. Support also came from the European Union’s Horizon Europe research and innovation programme under grant agreement No. 101097224 (REBECCA). Further financial assistance was provided by the Barcelona Zettascale Laboratory (reference REGAGE22e00058408992), which is financed by the Ministry for Digital Transformation and Public Services within the framework of the Resilience and Recovery Facility, and the European Union NextGenerationEU initiative.
We acknowledge the NMDB database (www.nmdb.eu), founded under the European Union’s FP7 programme (contract no. 213007), for supplying part of the neutron monitor data. Also, the Space Research Group (SRG-UAH) at the University of Alcala, Spain, for providing the CaLMa neutron monitor data. We further extend our gratitude to Darko Zivanović and Ferad Zyulkyarov for their collaboration and support.